# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ 6 Conclusion and limitation
This work presented the first rigorous guarantees on the smooth CE for several widely used learning algorithms. Our analysis proceeds in two stages: we first derive a uniform convergence bound for the smooth CE, and then upper bound the training smooth CE via functional gradients. We demonstrate that algorithms that can control the functional gradient simultaneously achieve a small smooth CE and misclassification rate. Despite these contributions, several limitations remain. First, for GBTs, our analysis requires a depth assumption that is stronger than what is typically necessary in practice, where shallow trees often yield well-calibrated predictions. Second, in the context of kernel boosting and NNs, our bounds are derived using a uniform bound over the post-processing function $h$, which may result in loose estimates. More refined analyses that better align with practical performance would be valuable. Moreover, our kernel boosting analysis relies on a strong margin assumption, which requires a well-separated data distribution. Since calibration is a rather weak notion compared with accuracy, extending the analysis to more realistic and weaker conditions is an important direction for future work. Finally, our analysis is limited to binary classification as the smooth CE is designed for this setting. Extending our analysis to the multiclass setting should be explored in future work.
@@ Appendix A Additional facts about proper losses
$\ell(p,y)$ $-y\ln p-(1-y)\ln(1-p)$ $(y-p)^{2}$
$\phi(v)$ $p\ln p+(1-p)\ln(1-p)$ $p(p-1)$
$\text{dual}(p)$ $\ln\left(\frac{p}{1-p}\right)$ $12p-$ 1
@@ B.1 True distance calibration and smooth CE
[…]
@@
[…]
@@ B.2 ECE and binning estimator
The binning estimator is commonly used to approximate this. Given a partition =$\mathcal{I}=\{I_{1},\dots,I_{m}\}$ of $[0,1]$ into intervals, we define:
@@
demonstrating robust soundness. However, ECE does not satisfy robust completeness; blasiok2023unify proved (Lemma 4.8) that for any $\epsilon\in{}^{+}$, there exists a distribution $\mathcal{D}$ such that $\mathrm{dCE}_{\mathcal{D}}(f)\leq\epsilon$ but $\mathrm{ECE}_{\mathcal{D}}(f)\geq 1/2-\epsilon$. This also highlights the discontinuity of ECE, which hinders its estimation from finite samples.
@@
$\mathrm{ECE}_{\mathcal{D}}(f)\leq\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})+(1+L)w_{\Gamma}(\mathcal{I})).$ (48)
@@
$\mathrm{ECE}_{\mathcal{D}}(f)=\mathbb{E}\left[|\mathbb{E}[Y|f=V]-V|\right]$ $=\sum_{i=1}^{m}\mathbb{E}[\mathbbm{1}_{V\in I_{i}}\cdot|\mathbb{E}[Y|f=V]-V|]$
$=\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f=V]-V||V\in I_{i}]$
@@
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}||\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]|$
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}|\mathbb{E}[Y|V\in I_{i}]-\mathbb{E}[V|V\in I_{i}]|$
@@
Finally, the first term can be bounded by using the Lipshitz continuity,
$\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}||\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]|$ (52)
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}||\mathbb{E}[Y|V]-\mathbb{E}[Y|V_{\mathcal{I}}]|+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}||\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]|$ (53)
@@
The term $\mathbb{E}||\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]|=0$ holds by the definition of conditional expectation; see also Lemma 1 in NEURIPS2024_futami. Thus, we have established Eq. (48).
Next, we consider taking the infimum over all partitions $\mathcal{I}$. By combining Claim 6.6 and Lemma 6.7 in blasiok2023unify, we obtain:
$\mathrm{ECE}_{\mathcal{D}}(f)\leq(1+2/\epsilon)\text{\text@underline{$\mathrm{dCE}_{\mathcal{D}}(f)$}}+(1+L)\epsilon$ (56)
@@ B.3 Other metrics
We numerically evaluate the behavior of MMCE with the Laplace kernel and smooth CE in Appendix H.
@@ C.1 Proof of Theorem 1
$-1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq|v_{i}-v_{j}|,\quad\forall i,j.$
@@
$|\mathrm{smCE}(f,\mathcal{D},S_{\mathrm{te}})-\mathrm{smCE}(f,\mathcal{D},S_{\mathrm{tr}})|\leq\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$ (62)
@@
Under the same setting as Theorem 1, we have
$\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\leq\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim D^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|+\sqrt{\frac{\log\frac{1}{\delta}}{n}}.$ (63)
@@
and thus the terms $\phi(f,z_{i})$ are dependent through $\omega^{*}_{i}$, which is a solution of the linear program in Eq. (60). Since there are $n$ dependent variables $\phi(f,z_{i})$, Hoeffding’s inequality, which requires independence, is not applicable. To establish Lemma 1, we instead employ McDiarmid’s inequality combined with a bounded difference argument; see the proof in Appendix C.2.
@@
To simplify the discussion, let us consider the case $n=2$. For a fixed dataset $S$ and function $f$, we have:
@@
Since the expectation of the exponential term is 0 from Eq. (82), we apply McDiarmid’s inequality:
@@
The proof of this theorem is provided below. Using this, we obtain
@@
Since $\mathcal{F}$ can be uncountable, we apply a discretization argument using the covering number. By the definition of the supremum, for any ϵ+superscriptitalic-$\epsilon^{\prime}\in\mathbb{R}^{+}$, there exists $f^{*}\in\mathcal{F}$ such that
@@
[…]
@@
We can use Eq. (106) for the uniform convergence bound, however, to get the refined dependency, we use the chaining technique; Let $\epsilon_{\ell}=2^{-\ell}$ for $\ell=0,1,2,\dots$. Let $\mathcal{F}_{\ell}$ be an $\epsilon_{\ell}$-cover of $\mathcal{F}$ with metric $L_{2}(S_{2n})$, and define $N_{\ell}=|\mathcal{F}_{\ell}|=N(\epsilon_{\ell},\mathcal{F},L_{2}(S_{2n}))$. We may set $\mathcal{F}_{0}=\{0\}$ at scale $\epsilon_{0}=1$.
@@
Similar to the derivation of Eqs. (105) and (106), regarding $\phi(f,Z_{i},Z^{\prime}_{i})$ as $f(Z_{i})$, we need to care that $2\epsilon$ appears. Therefore, given $\epsilon$-cover for $\mathcal{F}$,
@@
${\leq}\epsilon_{L}+\sum_{\ell=1}^{L}\sup_{f\in\mathcal{F}}\|f_{\ell}(f)-f_{\ell-1}(f)\|_{L_{2}(S_{n})}\sqrt{\frac{2\ln(N_{\ell}N_{\ell-1})}{n}}$ (109)
@@
[…]
Then for any $\epsilon>0$, we pick $L=\sup\{j:\epsilon_{L}>2\epsilon\}$. We then consider $\epsilon_{L+1}\leq 2\epsilon$, which implies $\epsilon_{L}\leq 4\epsilon$ and $\epsilon_{L}>2\epsilon$ implies $\epsilon_{L+1}>\epsilon$
@@
[…]
@@
$-1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|g(X_{i})-g(X_{j})|,\quad\forall i,j,$
@@
$-1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|\sigma^{-1}(v_{i})-\sigma^{-1}(v_{j})|,\quad\forall i,j.$
@@ C.2 Proof of Lemma 1
Therefore, we are required to estimate the constants $c_{i}$ in Eq. (115). To this end, consider replacing the $i$-th data point $z_{i}$ with $z^{\prime}_{i}$, and let the resulting dataset be denoted by $S^{\prime}_{n}=(z_{1},\dots,z_{i-1},z^{\prime}_{i},z_{i+1},\dots,z_{n})$. We define the following notation:
@@
where $\omega^{*}_{j}$ is the solution of Eq. (60) given $f$ and $S_{n}$. Similarly, define
@@
where $\omega^{{}^{\prime}*}_{j}$ is the solution of Eq. (60) given $f$ and $S^{\prime}_{n}$.
@@
$=\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{*}_{i}\leq\frac{1}{n},$ (124)
@@
$=\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{{}^{\prime}*}_{i}\geq-\frac{1}{n}.$ (128)
@@
$\left|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\right|\leq\frac{1}{n}.$ (129)
@@
Our goal is now to study the uniform stability of the quantity $\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$. By definition, for any ϵ+italic-$\epsilon\in\mathbb{R}^{+}$, there exists $g\in\mathcal{F}$ such that
@@
We first consider the case where the replaced data point is from the training set $S_{\mathrm{te}}=\{\tilde{z}_{i}\}_{i=1}^{n}$. Let the perturbed dataset be
@@
Then, using the triangle inequality and Eq. (130), we have:
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\leq|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\leq\frac{1}{n},$ (132)
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\geq-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\geq-\frac{1}{n}.$ (133)
@@
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|\right|\leq\frac{1}{n}.$ (134)
A similar analysis applies when the replaced data point is from the test set $S_{\mathrm{tr}}=\{z_{i}\}_{i=1}^{n}$. Let the perturbed test set be
@@
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}}^{\prime})|\right|\leq\frac{1}{n}.$ (135)
@@
$c_{i}=\sup_{\{z_{j}\}_{j=1}^{2n},z^{\prime}_{i}}\left|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})-\left(\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})\right)\right|\leq\frac{1}{n},$ (136)
@@
Applying McDiarmid’s inequality (Eq. (117)) with these coefficients yields:
$\Pr\left(\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|-\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim\mathcal{D}^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\geq\epsilon\right)\leq\exp(-n\epsilon^{2}).$ (137)

### v4

@@ 6 Conclusion and limitation
This work presented the first rigorous guarantees on the smooth CE for several widely used learning algorithms. Our analysis proceeds in two stages: we first derive a uniform convergence bound for the smooth CE, and then upper-bound the training smooth CE via functional gradients. We demonstrate that algorithms that can control the functional gradient simultaneously achieve a small smooth CE and misclassification rate. Despite these contributions, several limitations remain. First, our bounds are derived using a uniform bound over the post-processing function $h$, not the Lipschitz function class, which may result in loose estimates. More refined analyses that better align with practical performance would be valuable. Moreover, our analysis relies on a strong margin assumption, which requires a well-separated data distribution. Since calibration is a rather weak notion compared with accuracy, extending the analysis to more realistic and weaker conditions is an important direction for future work. Finally, our analysis is limited to binary classification as the smooth CE is designed for this setting. Extending our analysis to the multiclass setting should be explored in future work.
@@ Appendix A Additional facts about proper losses
$\ell(p,y)$ $y\ln p-(1-y)\ln(1-p)$ $(y-p)^{2}$
$\phi(v)$ $p\ln p+(1-p)\ln(1-p)$ p(p1)p(p-1)
$\text{dual}(p)$ $\ln\left(\frac{p}{1-p}\right)$ 2p12p-1
@@ B.1 True distance calibration and smooth CE
We say that a prediction-label distribution $\Gamma$ over $0,1]\times\{0,1\}$ is perfectly calibrated if $\mathbb{E}_{(v,y)\sim\Gamma}[y|v]=v$. Moreover, given $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$, we say that $f$ is perfectly calibrated with respect to $\mathcal{D}$ if $\mathcal{D}_{f}$ is perfectly calibrated.
@@
For $c>0$, a calibration metric $\mu$ is said to satisfy $c$-robust completeness if there exists a constant $a\geq 0$ such that for every distributionDD on $\mathcal{X}\times\{0,1\}$, and predictor $f\in\mathcal{F}$
@@ B.2 ECE and binning estimator
The binning estimator is commonly used to approximate this. Given a partition $\mathcal{I}=\{I_{1},\dots,I_{m}\}$ of $[0,1]$ into intervals, we define:
@@
demonstrating robust soundness. However, ECE does not satisfy robust completeness; blasiok2023unify proved (Lemma 4.8) that for any $\epsilon\in\real^{+}$, there exists a distribution $\mathcal{D}$ such that $\mathrm{dCE}_{\mathcal{D}}(f)\leq\epsilon$ but $\mathrm{ECE}_{\mathcal{D}}(f)\geq 1/2-\epsilon$. This also highlights the discontinuity of ECE, which hinders its estimation from finite samples.
@@
$\mathrm{ECE}_{\mathcal{D}}(f)\leq\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})+(1+L)w_{\Gamma}(\mathcal{I}).$ (49)
@@
$\mathrm{ECE}_{\mathcal{D}}(f)=\mathbb{E}\left[|\mathbb{E}[Y|f=V]-V|\right]$ $=\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f=V]-V||V\in I_{i}]$
@@
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V\in I_{i}]-\mathbb{E}[V|V\in I_{i}]||V\in I_{i}]$
@@
Finally, the first term can be bounded by using the Lipschitz continuity,
$\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$ (53)
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V_{\mathcal{I}}]||V\in I_{i}]$ (54)
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$ (55)
@@
The term $\mathbb{E}[|\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]=0$ holds by the definition of conditional expectation. Thus, we have established Eq. equation 49.
Next, we consider taking the infimum over all partitions $\mathcal{I}$. By combining Claim 6.6 and Lemma 6.7 in blasiok2023unify and Eq. equation 49, we obtain:
[…]
@@ B.3 Other metrics
We numerically evaluate the behavior of MMCE with the Laplace kernel and smooth CE in Appendix I.
@@ C.1 Proof of Theorem 1
Before the formal proof, we provide a proof sketch to highlight the key differences from standard generalization bounds.
To address this, we introduce $\{\sigma_{i}\}$s in a way that preserves the structure of the linear convex formulation of smooth CE. Since the resulting bound does not take the form of standard Rademacher complexity, we discretize the hypothesis class $\mathcal{F}$ and derive a covering number bound.
@@
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq|v_{i}-v_{j}|,\quad\forall i,j.$
@@
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$ (64)
@@
Under the same setting as Theorem 1, with probability $1-\delta$, we have
$\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\leq\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim D^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|+2\sqrt{\frac{\log\frac{1}{\delta}}{n}}.$ (65)
The proof of this lemma is provided in Appedix C.2.
@@
and thus the terms $\phi(f,z_{i})$ are dependent through $\omega^{*}_{i}$, which is a solution of the linear program in Eq. equation 62. Since there are $n$ dependent variables $\phi(f,z_{i})$, Hoeffding’s inequality, which requires independence, is not applicable. To establish Lemma 1, we instead employ McDiarmid’s inequality combined with a bounded difference argument; see the proof in Appendix C.2.
@@
To simplify the discussion, let us consider the case $n=2$. For a fixed dataset $S$ and function $f$, Eq. equation 67 under the expression of Eq. equation 68 is written as:
@@
Since the expectation of the exponential term is 0 from Eq. equation 84, we apply McDiarmid’s inequality:
@@
The proof of this theorem is provided in Appendix C.3. Using this, we obtain
@@
Since $\mathcal{F}$ can be uncountable, we apply a discretization argument using the covering number. By the definition of the supremum, for any $\epsilon^{\prime}\in\mathbb{R}^{+}$, there exists $f^{*}\in\mathcal{F}$ such that
@@
$\leq\text{(first term)}+\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\tilde{f}(x_{i})-f^{*}(x_{i})|^{2}}+\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\tilde{f}(x^{\prime}_{i})-f^{*}(x^{\prime}_{i})|^{2}}$ (105)
$\leq\text{(first term)}+2\epsilon,$ (106)
@@
We can use Eq. equation 108 for the uniform convergence bound, however, to get the refined dependency, we use the chaining technique; Let $\epsilon_{\ell}=2^{-\ell}$ for $\ell=0,1,2,\dots$. Let $\mathcal{F}_{\ell}$ be an $\epsilon_{\ell}$-cover of $\mathcal{F}$ with metric $L_{2}(S_{2n})$, and define $N_{\ell}=|\mathcal{F}_{\ell}|=N(\epsilon_{\ell},\mathcal{F},L_{2}(S_{2n}))$. We may set $\mathcal{F}_{0}=\{0\}$ at scale $\epsilon_{0}=1$.
@@
Similar to the derivation of Eqs. equation 107 and equation 108, regarding $\phi(f,Z_{i},Z^{\prime}_{i})$ as $f(Z_{i})$, we need to care that $2\epsilon$ appears. Therefore, given $\epsilon$-cover for $\mathcal{F}$,
@@
${\leq}2\epsilon_{L}+\sum_{\ell=1}^{L}\sup_{f\in\mathcal{F}}\|f_{\ell}(f)-f_{\ell-1}(f)\|_{L_{2}(S_{n})}\sqrt{\frac{2\ln(N_{\ell}N_{\ell-1})}{n}}$ (111)
@@
$\leq 2\epsilon_{L}+24\int_{\epsilon_{L+1}}^{\epsilon_{0}}\frac{\sqrt{\ln N(\epsilon^{\prime},\mathcal{F},L_{2}(S_{n}))}}{\sqrt{n}}d\epsilon^{\prime}.$ (114)
Then for any $\epsilon>0$, we pick $L=\sup\{j:\epsilon_{j}>2\epsilon\}$. This simple $\epsilon_{L+1}\leq 2\epsilon$, thus $\epsilon_{L}\leq 4\epsilon$ holds. By definition of $L$, $\epsilon_{L}>2\epsilon$ and this implies $\epsilon_{L+1}>\epsilon$
@@
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq\inf_{\epsilon\geq 0}\left[8\epsilon+24\int_{\epsilon}^{1}\frac{\sqrt{\ln N(\epsilon^{\prime},\mathcal{F},L_{2}(S_{n}))}}{\sqrt{n}}d\epsilon^{\prime}\right].$ (115)
@@
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|g(X_{i})-g(X_{j})|,\quad\forall i,j,$
@@
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|\sigma^{-1}(v_{i})-\sigma^{-1}(v_{j})|,\quad\forall i,j.$
@@ C.2 Proof of Lemma 1
Therefore, we are required to estimate the constants $c_{i}$ in Eq. equation 117. To this end, consider replacing the $i$-th data point $z_{i}$ with $z^{\prime}_{i}$, and let the resulting dataset be denoted by $S^{\prime}_{n}=(z_{1},\dots,z_{i-1},z^{\prime}_{i},z_{i+1},\dots,z_{n})$. We define the following notation:
@@
where $\omega^{*}_{j}$ is the solution of Eq. equation 62 given $f$ and $S_{n}$. Similarly, define
@@
where $\omega^{{}^{\prime}*}_{j}$ is the solution of Eq. equation 62 given $f$ and $S^{\prime}_{n}$.
@@
$=\frac{1}{n}(y_{i}-y^{\prime}_{i})\omega^{*}_{i}+\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{*}_{i}\leq\frac{2}{n},$ (126)
@@
$=\frac{1}{n}(y^{\prime}_{i}-y_{i})\omega^{*}_{i}+\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{{}^{\prime}*}_{i}\geq-\frac{2}{n}.$ (130)
@@
$\left|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\right|\leq\frac{2}{n}.$ (131)
@@
Our goal is now to study the uniform stability of the quantity $\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$. By definition, for any $\epsilon\in\mathbb{R}^{+}$, there exists $g\in\mathcal{F}$ such that
@@
We first consider the case where the replaced data point is from the test set $S_{\mathrm{te}}=\{\tilde{z}_{i}\}_{i=1}^{n}$. Let the perturbed dataset be
@@
Then, using the triangle inequality and Eq. equation 132, we have:
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\leq|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\leq\frac{2}{n},$ (134)
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\geq-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\geq-\frac{2}{n}.$ (135)
@@
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|\right|\leq\frac{2}{n}.$ (136)
A similar analysis applies when the replaced data point is from the training set $S_{\mathrm{tr}}=\{z_{i}\}_{i=1}^{n}$. Let the perturbed test set be
@@
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}}^{\prime})|\right|\leq\frac{2}{n}.$ (137)
@@
$c_{i}=\sup_{\{z_{j}\}_{j=1}^{2n},z^{\prime}_{i}}\left|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})-\left(\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})\right)\right|\leq\frac{2}{n},$ (138)
@@
Applying McDiarmid’s inequality (Eq. equation 119) with these coefficients yields:
$\Pr\left(\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|-\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim\mathcal{D}^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\geq\epsilon\right)\leq\exp(-n\epsilon^{2}/4).$ (139)